# Data-driven model validation for neutrino-nucleus cross section measurements

### v1

@@ II.1 A Priori Information About Event Generators
This motivates the need for data-driven inputs to inform the allowed parameters for these models and reduce their uncertainties. Experiments often supplement generators with tailored “tunes” to better represent their data [18, 19, 20]. The success of this strategy in fulfilling the requirement of generating complete final-state particle kinematics in neutrino-nucleus interactions has been demonstrated with the consistent $|\Delta m^{2}_{\mathrm{atm}}|$ values extracted from the accelerator neutrino oscillation experiments [1, 2] and reactor antineutrino oscillation experiments [21, 22], in which the inverse $\beta$ decay process allows for a simpler and more precise reconstruction of neutrino energy. Nevertheless, because of their hybrid nature, event generators tend to better describe inclusive processes or phase spaces where sufficient data were accumulated. For other exclusive processes, which require a more detailed description of the hadronic final states, event generators are more likely to fall short.
This places experiments in the following situation. On one hand, it is unlikely the parameters available in current event generators are sufficient in describing interaction modes in the complete phase space. On the other hand, the conservative uncertainties assigned on these parameters partially mitigate this shortcoming but often lead to large systematic uncertainties. As such, cross section measurements that treat the reliance on event generator with care remain essential in stimulating improvements to simulation that will enable the desired level of precision in current and future neutrino experiments.
@@ II.2 Sources of Model Dependence
From Eq. 1, it is clear that cross section extraction depends on a model prediction for the neutrino flux, the rate of background events, the selection efficiency, and detector effects that result in the imperfect reconstruction of kinematic quantities accounted for in the smearing matrix. This dependence on the overall model means that mis-modeling in the phase space relevant to the cross section extraction can introduce bias into the measurement. For measurements that must contend with low efficiencies, purities, or substantial detector effects, this phase space may extend beyond the measured one into background channels, where mismodeling could lead to an incorrect background correction, or into channels and observables adjacent to the measurement where mismodeling could suggest an incorrect efficiency estimation. The fundamental challenge here is that one does not know if the model used in the extraction can describe nature. Moreover, when the overall model is unable to describe nature, it is usually unclear if the discrepancy is due to detector, cross section, or flux effects and may even be due to a combination of a variety of sources of mis-modeling. In these cases, there may be a need for additional uncertainties beyond those inside the model, and therefore a form of model validation is required to verify that the existing uncertainties provide sufficient coverage of the data.
@@
Visible kinematic variables do not entirely avoid model dependence either. Whenever a measurement relies upon mapping from reconstructed quantities to true quantities this mapping must be estimated through a model, motivating the need for model validation in these cases. In particular, the mapping for quantities like the available energy $E_{\mathrm{avail}}$ [23, 24, 25, 26, 12], often defined as the sum of reconstructable energy deposited by visible particles, serves as an alternative to the energy transfer but has a strong dependence on the accurate simulation of particles that deposit energy in the detector. Since different modeling and reconstruction failures may be present in different final states, the mapping from reconstructed to true $E_{\mathrm{avail}}$ requires a robust description of the complete contents of the hadronic final state, thereby making it susceptible to model dependence.
@@ II.3 Real versus Nominal Flux
The extent of this model dependence depends on subtleties in the treatment of the flux and its related uncertainties. In the literature, this has been described as whether the measurement is averaged over an assumed well defined nominal neutrino flux spectrum, or averaged over the unknown real neutrino flux spectrum impingent on the detector [28]. When extracting cross sections in the real flux, unfolding amounts to correcting for detector effects with a presumed cross section and flux model. When extracting cross sections in a nominal flux, one corrects for detector effects but then also extrapolates from the (unknown) real flux to the nominal flux in the unfolding. This distinction is subtle but leads to important differences in the treatment of flux uncertainties. These uncertainties can be quite prominent in the context of accelerator-based neutrino experiments, which often have 10% or higher uncertainties on their flux predictions.
@@
If an analysis extracts cross sections in the nominal neutrino flux spectrum, model-dependence arises when the mapping between neutrino energy and the visible kinematic variable is used to extrapolate from the real to the nominal neutrino spectrum. In this approach, flux uncertainties are estimated by varying the assumed real flux, but not the well-defined nominal flux, when determining the impact of flux effects. Referring to Eq. 1, this amounts to keeping $\Phi$ constant while re-evaluating $n_{a}$ based on the real flux in each flux systematic universe. This approach allows the covariance matrix to include the uncertainties of extrapolating the data from the unknown real neutrino flux to the nominal flux and corresponding $\Phi$ that the results are averaged over. In this case, comparisons to external theory or event generator predictions are straightforward because there is no need for future analyzers of the data to utilize additional flux uncertainties. Flux systematics are entirely accounted for by the experimentalist when estimating uncertainties for the cross section extraction. In this case, data-driven model validation can serve as a mechanism to quantify the bias introduced in the extrapolation from real to nominal flux and help avoid under-estimating or over-estimating flux uncertainties.
@@
As demonstrated by this toy study, the requirements of reporting and using neutrino flux uncertainties and their correlations with cross section results make it very difficult, if not impossible, for theorists to properly compare their predictions with cross sections extracted using the real neutrino spectrum. Indeed, this may be related to some of the issues encountered in recent efforts to tune event generators to experimental data [18, 30, 31]. Because of this, we advocate for extracting cross sections in the nominal neutrino flux spectrum, which allows the results to be reported in a single well-defined flux. With this method, all flux uncertainties are included in the covariance matrix obtained in the unfolding. However, this method also introduces additional model dependence on the experimental side associated with the mapping between the neutrino energy spectrum and the measured kinematic variables. We thus advocate for data-driven model validation when extracting cross sections in any physics variable at the nominal neutrino flux spectrum to ensure that the uncertainties associated with extrapolating from the true to the nominal neutrino flux spectrum are sufficient.
@@ III.1.2 $\chi^{2}$ Decompositions
In an overall GoF test, it is possible that conservative uncertainties have hidden a significant discrepancy between data and model in some bins that may bias the cross section extraction in select regions of phase space. However, it is challenging to further evaluate the GoF in select regions of phase space because there are generally strong bin-to-bin correlations in the reconstructed distributions. To address this, the overall test statistic can be decomposed by diagonalizing the covariance matrix, thereby making all bins uncorrelated. In this transformed basis, the absence of correlations allows for a rigorous quantification of the GoF between the data and model prediction within a single bin given the uncertainties of the diagonalized covariance matrix. This allows one to test the local GoF of distributions in the decomposed space.
@@
An example of utilizing the $\chi^{2}$ decomposition is illustrated in Fig. 2 on MicroBooNE data. The distribution of interest in this figure is the $\nu_{\mu}$ selection from [34] binned as a function of the reconstructed hadronic energy, which Fig. 2(b) shows in reconstructed space. Figure 2(c) then shows the the $\chi^{2}$ decomposition and includes both the resulting distribution of $\epsilon_{i}$ values in the decomposition space and the matrix used to transform from the reconstructed space to the decomposition space. This matrix, though generally challenging to interpret, can provide some insight into the mapping between the two spaces and the types of discrepancies that would result in significant tension in individual $\epsilon_{i}$. Taking the first decomposition bin as an example, we see that this bin receives a negative contribution from the first three reconstructed bins and a positive contribution from all higher energy bins. As such, a migration of events from these first three low energy bins into the higher energy bins would likely cause a large discrepancy in this decomposition bin which would likewise result in a large $\epsilon_{i}$ indicative of significant mis-modeling.
@@ III.1.3 Conditional Constraints
The procedure described in Eqs. 10 and 11 allows a GoF test to be performed on $X$ after the constraints from $Y$. This allows the simultaneous examination of the correlated modeling of $X$ and $Y$, which provides additional information about the compatibility between the model and data. Typically, the constrained and constraining channels have highly correlated model predictions due to detector and physics effects that impact both sets of distributions. The model’s description of the correlations is dictated by the modeling of these effects. The reduction in uncertainties for the prediction on the constrained distribution provides additional sensitivity by providing a means of exploring the modeling of these correlations. The modeling of $X$, $Y$, and the relationship between $X$ and $Y$ must all be sufficient for the model to pass a constrained validation test.
@@
An example of using the conditional constraint can be seen in Fig. 2. In this example, $Y$ is defined to be the distributions of the reconstructed muon kinematics, namely, the reconstructed muon energy, $E_{\mu}^{\mathrm{rec}}$, and the muon angle, $\cos\theta_{\mu}^{\mathrm{rec}}$, where $\theta_{\mu}^{\mathrm{rec}}$ is defined as the angle the muon scatters with respect to the incoming neutrino beam. These distributions are shown in Fig. 2(a). Events fully contained (FC) and partially contained (PC) within the detector are placed into separate bins to better separate these two classes of events, which may be sensitive to different forms of mis-modeling. Distribution $X$ s chosen to be that of reconstructed hadronic, $E_{\mathrm{had}}^{\mathrm{rec}}$, for events partially contained (PC) within the detector. This distribution is shown in Fig. 2(b). Note that in this case, $X$ and $Y$ contain the same events, and therefore statistical correlations are present in the covariance matrix. The MicroBooNE Monte Carlo (MC) prediction before constraint is shown in red and the prediction after constraint is shown in blue. The shift between the red and blue predictions is dictated by the observation in data for the muon kinematics and by the model’s predicted relationship between the muon kinematics and the reconstructed hadronic energy. The bands surrounding the prediction, which show the uncertainties before and after constraint, illustrate the large reduction in uncertainties and enhanced sensitivity to mis-modeling that is obtainable with the conditional constraint. In particular, this test is expected to provide sensitivity to the modeling of the missing hadronic energy as is described in more detail in Sec. III.2. Other examples of similar constraint tests can be found in [10, 11, 12].
@@ III.2 Developing a Model Validation Procedure
However, this also brings about its own set of downsides. A combination of flux, detector and cross section effects could conspire in such a way to cancel out, thereby making reconstructed space tests less sensitive to mis-modeling that would still bias the extracted cross sections. The potential for such a situation is investigated in Sec IV.2.1 and should be kept in mind when deciding which tests to include in the model validation procedure. In addition, it is important to note that the procedures we outline are more applicable when the background and signal models are both well defined and relatively well understood up to their uncertainties. If a signal significantly deviates from existing models or is completely unknown then these techniques cannot be properly applied as validation. In that case, other techniques should be employed to validate the background modeling and the signal efficiency within uncertainties before extracting a cross section.
@@
When developing a model validation procedure for a desired cross section extraction, one must make sure to design a set of tests that are sufficiently sensitive to mis-modeling in the phase space relevant for the cross section extraction. For example, validating only the overall event rate will not be sufficient for a differential cross section measurement. Such a test averages over the entire reconstructed phase space and will thus be insensitive to any mis-modeling related to the shape of the distribution that could bias the extraction of a differential cross section. Selecting an appropriately sensitive set of tests represents the key to a well-designed model validation procedure. To achieve this, one should consider what forms of mismodeling may be capable of introducing bias into their cross section result and choose tests that address these possibilities. Mismodeling in distributions irrelevant to the extraction do not need to be examined. It may be useful to use FDSs to evaluate the ability of the selected set of tests to detect relevant mis-modeling before it begins to bias the extracted cross section results beyond stated uncertainties. A case study containing this type of FDS is presented in Sec. IV. This study mirrors the analysis done in [10] and thus also provides an example of what a full suite of tests used in a data-driven validation may look like. However, such studies are not mathematical proofs that the intended validation procedure constructed for the given analysis is guaranteed to detect relevant mis-modeling. As such, it is pertinent to probe the model from a variety of different angles in order to maximize the probability of detecting problematic forms of mis-modeling that would bias the cross section results beyond their uncertainties.
@@
The more challenging a variable or channel is to reconstruct or model, the more stringently it should be examined in the validation. For visible kinematic variables in more inclusive cross section measurements, a direct data to prediction comparison over the reconstructed distribution used in the unfolding is likely sufficient. However, more challenging distributions to model or reconstruct require additional validation. Examples include distributions containing events that are not fully contained withing the detector volume, distributions used to map to quantities that cannot be fully reconstructed, or a more exclusive channel with kinematics cuts that are near detector thresholds.
@@
One particularly prominent example of this is the reconstruction of low energy protons. Low energy nucleons are known to be modeled significantly differently among event generators [39, 40] thus leading to very different predicted selection efficiencies for protons near the detection threshold [27]. This can have a large impact on measurements related to the proton’s energy, such as the kinetic energy of the leading proton or transverse kinematic imbalance variables [41, 42, 43], which are projections of the lepton and proton momentum onto the plane perpendicular to the neutrino direction. In such a case, a conditional constraint from a related visible kinematic variable less susceptible to the threshold may be useful. Examples of this include the energy and angle of the outgoing lepton. These constraints will reduce the allowed parameter space through correlations due to shared detector, flux, and cross section modeling thereby allowing for a more thorough investigation of the modeling of the proton kinematics near the detection threshold. Including an additional constraint from a closely related channel, such as one without protons, can provide additional information on the modeling of backgrounds or signal events that did not pass selection cuts and may also be useful in probing for relevant mis-modeling. This can be seen as similar to studying side-bands, which are often used to evaluate the modeling of backgrounds, but intends to evaluate the modeling used in the cross section extraction more directly.
@@
Through the use of $E_{\ell}^{\mathrm{rec}}$ as a constraint, the simultaneous measurement of $E_{\mathrm{had}}^{\mathrm{rec}}$ and $E_{\ell}^{\mathrm{rec}}$ is able to validate the predicted relationship between these distributions. This is achieved by using $E_{\ell}^{\mathrm{rec}}$ as channel $Y$ and $E_{\mathrm{had}}^{\mathrm{rec}}$ as channel $X$ in Eq. 10:
@@
where $\Sigma^{E_{\ell}^{\mathrm{rec}}E_{\ell}^{\mathrm{rec}}}$ describes the uncertainties on $E_{\ell}^{\mathrm{rec}}$ and $\Sigma^{E_{\ell}^{\mathrm{rec}}E_{\mathrm{had}}^{\mathrm{rec}}}$ describes the correlations between $E_{\ell}^{\mathrm{rec}}$ and $E_{\mathrm{had}}^{\mathrm{rec}}$. The same is done for Eq. 11 to obtain the reduced uncertainty band, $\Sigma^{E_{\mathrm{had}}^{\mathrm{rec}}E_{\mathrm{had}}^{\mathrm{rec}},\text{const.}}$, for the $E_{\mathrm{had}}^{\mathrm{rec}}$ prediction. From Eq. 13, we see that the modeling of $E_{\ell}^{\mathrm{rec}}$, $E_{\mathrm{had}}^{\mathrm{rec}}$, and the correlations between them all play a role in obtaining the posterior prediction. Following Eq. 4, these can then be simultaneously examined with a GoF test on the constrained $E_{\mathrm{had}}^{\mathrm{rec}}$ prediction:

### v2

@@ II.1 A Priori Information About Event Generators
This motivates the need for data-driven inputs to inform the allowed parameters for these models and reduce their uncertainties. Experiments often supplement generators with tailored “tunes” to better represent their data [20, 21, 22, 23, 24, 25]. The success of this strategy in fulfilling the requirement of generating complete final-state particle kinematics in neutrino-nucleus interactions has been demonstrated with the consistent $|\Delta m^{2}_{\mathrm{atm}}|$ values extracted from the accelerator neutrino oscillation experiments [1, 2] and reactor antineutrino oscillation experiments [26, 27], in which the inverse $\beta$ decay process allows for a simpler and more precise reconstruction of neutrino energy. Nevertheless, because of their hybrid nature, event generators tend to better describe inclusive processes or phase spaces where sufficient data were accumulated. For other exclusive processes, which require a more detailed description of the hadronic final states, event generators are more likely to fall short.
This places experiments in the following situation. On one hand, it is unlikely the parameters available in current event generators are sufficient in describing interaction modes in the complete phase space. On the other hand, the conservative uncertainties assigned on these parameters partially mitigate this shortcoming but often lead to large systematic uncertainties. As such, cross section measurements that treat the reliance on event generators with care remain essential in stimulating improvements to simulation that will enable the desired level of precision in current and future neutrino experiments.
@@ II.2 Sources of Model Dependence
From Eq. 1, it is clear that cross section extraction depends on a model prediction for the neutrino flux, the rate of signal and background events, the selection efficiency, and detector effects that result in the imperfect reconstruction of kinematic quantities accounted for in the smearing matrix. This dependence on the overall model means that mis-modeling in the phase space relevant to the cross section extraction can introduce bias into the measurement. For measurements that must contend with low efficiencies, purities, or substantial detector effects, this phase space may extend beyond the measured one into background channels, where mis-modeling could lead to an incorrect background correction, or into channels and observables adjacent to the measurement where mis-modeling could suggest an incorrect efficiency estimation. The fundamental challenge here is that one does not know if the model used in the extraction can describe nature. Moreover, when the overall model is unable to describe nature, it is usually unclear if the discrepancy is due to detector, cross section, or flux effects and may even be due to a combination of a variety of sources of mis-modeling. In these cases, there may be a need for additional uncertainties beyond those inside the model, and therefore a form of model validation is required to verify that the existing uncertainties provide sufficient coverage of the data.
@@
Visible kinematic variables do not entirely avoid model dependence either. Whenever a measurement relies upon mapping from reconstructed quantities to true quantities this mapping must be estimated through a model, motivating the need for model validation in these cases. In particular, the mapping for quantities like the available energy $E_{\mathrm{avail}}$ [28, 29, 30, 31, 12, 32, 18], often defined as the sum of reconstructable energy deposited by visible particles, serves as an alternative to the energy transfer but has a strong dependence on the accurate simulation of particles that deposit energy in the detector. Since different modeling and reconstruction failures may be present in different final states, the mapping from reconstructed to true $E_{\mathrm{avail}}$ requires a robust description of the complete contents of the hadronic final state, thereby making it susceptible to model dependence.
@@ II.3 Real versus Nominal Flux
The extent of this model dependence depends on subtleties in the treatment of the flux and its related uncertainties. In the literature, this has been described as whether the measurement is averaged over an assumed well defined nominal neutrino flux spectrum, or averaged over the unknown real neutrino flux spectrum impingent on the detector [33]. When extracting cross sections in the real flux, unfolding amounts to correcting for detector effects with a presumed cross section and flux model. When extracting cross sections in a nominal flux, one corrects for detector effects but then also extrapolates from the (unknown) real flux to the nominal flux in the unfolding. This distinction is subtle but leads to important differences in the treatment of flux uncertainties. These uncertainties can be quite prominent in the context of accelerator-based neutrino cross section measurements, for which flux uncertainties are typically 5-10%, though can be even larger in instances where the flux is poorly known.
@@
If an analysis extracts cross sections in the nominal neutrino flux spectrum, model-dependence arises when the mapping between neutrino energy and the visible kinematic variable is used to extrapolate from the real to the nominal neutrino spectrum. In this approach, flux uncertainties are estimated by varying the assumed real flux, but not the well-defined nominal flux, when determining the impact of flux effects. Referring to Eq. 1, this amounts to keeping $\Phi$ constant while re-evaluating $n_{a}$ based on the real flux in each flux systematic universe. This approach allows the covariance matrix to include the uncertainties of extrapolating the data from the unknown real neutrino flux to the nominal flux and corresponding $\Phi$ that the results are averaged over. In this case, comparisons to external theory or event generator predictions are straightforward because there is no need for future analyzers of the data to utilize additional flux uncertainties. Flux systematics are entirely accounted for by the experimentalist when estimating uncertainties for the cross section extraction. In this case, data-driven model validation can serve as a mechanism to quantify the bias introduced in the extrapolation from real to nominal flux and help avoid under-estimating or over-estimating flux uncertainties. This is similar to the utility of data-driven model validation when extracting derived quantities dependent on the neutrino flux spectrum, such as the cross section as a function of the neutrino energy or energy transferred to the nucleus, where it likewise serves to validate the mapping between reconstructed quantities and the incoming neutrino energy spectrum. The dependence on the model is greater for these derived quantities, hence a greater need for careful validation, but one should still be cautious in entirely ignoring this model dependence in the context of visible kinematic variables.
@@
As demonstrated by this toy study, the requirements of reporting and using neutrino flux uncertainties and their correlations with cross section results make it very difficult, if not impossible, for theorists to properly compare their predictions with cross sections extracted using the real neutrino spectrum. Indeed, this may be related to some of the issues encountered in recent efforts to tune event generators to experimental data [20, 35, 24]. Because of this, we advocate for extracting cross sections in the nominal neutrino flux spectrum, which allows the results to be reported in a single well-defined flux. With this method, all flux uncertainties are included in the covariance matrix obtained in the unfolding. However, this method also introduces additional model dependence on the experimental side associated with the mapping between the neutrino energy spectrum and the measured kinematic variables. We thus advocate for data-driven model validation when extracting cross sections in any physics variable at the nominal neutrino flux spectrum. These methods are well suited to examine the entire uncertainty budget and thus help to ensure that the uncertainties associated with extrapolating from the true to the nominal neutrino flux spectrum are sufficient.
@@ III.1.2 $\chi^{2}$ Decompositions
In an overall GoF test, it is possible that conservative uncertainties have hidden a significant discrepancy between data and model in some bins that may bias the cross section extraction in select regions of phase space. However, it is challenging to further evaluate the GoF in select regions of phase space because there are generally strong bin-to-bin correlations in the reconstructed distributions. To address this, the overall test statistic can be decomposed by diagonalizing the covariance matrix, thereby making all bins uncorrelated. In this transformed basis, the absence of correlations allows for a rigorous quantification of the GoF between the data and model prediction within a single bin given the uncertainties of the diagonalized covariance matrix. This allows one to test the local GoF of distributions in the decomposed space. Though it is challenging to draw physics conclusion from this decomposed space, the technique is a useful statistical tool in evaluating the compatibility between data and prediction when significant correlations in reconstructed distributions may render the overall $\chi^{2}$ test statistic insensitive to local discrepancies.
@@
An example of utilizing the $\chi^{2}$ decomposition is illustrated in Fig. 2 on MicroBooNE data. The distribution of interest in this figure is the $\nu_{\mu}$ selection from [38] binned as a function of the reconstructed hadronic energy, which Fig. 2(b) shows in reconstructed space. Figure 2(c) then shows the the $\chi^{2}$ decomposition and includes both the resulting distribution of $\epsilon_{i}$ values in the decomposition space and the matrix used to transform from the reconstructed space to the decomposition space. This matrix, though generally challenging to interpret, can provide some insight into the mapping between the two spaces and the types of discrepancies that would result in significant tension in individual $\epsilon_{i}$. Taking the first decomposition bin as an example, we see that this bin receives a negative contribution from the first three reconstructed bins and a positive contribution from all higher energy bins. As such, an effect that modifies the event rate differently for the first three low energy bins and the higher energy bins would likely cause a large discrepancy in this decomposition bin which would likewise result in a large $\epsilon_{i}$ indicative of significant mis-modeling.
@@ III.1.3 Conditional Constraints
The amount by which the conditional mean differs from the prior central value prediction will be dictated by the three factors in the second term on the right of Eq. 10. Smaller uncertainties on the constraining distribution results in a more powerful constraint and a larger update to the prediction, this enters through $(\Sigma^{YY})^{-1}$. The correlations between the two channels enters through $\Sigma^{XY}$, with a higher level of correlation increasing the size of the update to the prediction. The amount the prediction in the constraining channels differs from the data observation enters through $(n^{Y}-\mu^{Y})$, with a larger deviation from the prediction producing a larger update to the prediction in the constrained channel. The reduction in uncertainties on the constrained distribution is described by the second term on the right in Eq. 11. Likewise, this term has factors that depend on the size of the uncertainties on the constraining distribution and the size of the correlation between the two channels, namely $(\Sigma^{YY})^{-1}$ and $\Sigma^{XY}$, respectively. Smaller uncertainties on the constraining channel and larger correlations will increase the resulting reduction in uncertainties on the constrained channel. However, unlike the second term in Eq. 10, the second term in Eq. 11 does not depend on the data observation in the constraining channel; the amount of “additional information” provided by the constraint remains constant regardless of the data observation.
@@
The procedure described in Eqs. 10 and 11 allows a GoF test to be performed on $X$ after the constraints from $Y$. This allows the simultaneous examination of the correlated modeling of $X$ and $Y$, which provides additional information about the compatibility between the model and data. Typically, the constrained and constraining channels have highly correlated model predictions due to detector and physics effects that impact both sets of distributions. The model’s description of the correlations is dictated by the modeling of these effects. The reduction in uncertainties for the prediction on the constrained distribution provides additional sensitivity by providing a means of exploring the modeling of these correlations. This direct incorporation of the correlations into the updated central value prediction, uncertainties, and subsequent GoF tests make it straightforward to perform a rigorous examination of the consistency of the modeling across channels and observables, which is a tangible advantage of this method. The modeling of $X$, $Y$, and the relationship between $X$ and $Y$ must all be sufficient for the model to pass a constrained validation test.
@@
An example of using the conditional constraint can be seen in Fig. 2. In this example, $Y$ is defined to be the distributions of the reconstructed muon kinematics, namely, the reconstructed muon energy, $E_{\mu}^{\mathrm{rec}}$, and the muon angle, $\cos\theta_{\mu}^{\mathrm{rec}}$, where $\theta_{\mu}^{\mathrm{rec}}$ is defined as the angle the muon scatters with respect to the incoming neutrino beam. These distributions are shown in Fig. 2(a). Events fully contained (FC) and partially contained (PC) within the detector are placed into separate bins to better separate these two classes of events, which may be sensitive to different forms of mis-modeling. Distribution $X$ is chosen to be that of reconstructed hadronic energy, $E_{\mathrm{had}}^{\mathrm{rec}}$, for events partially contained (PC) within the detector. This distribution is shown in Fig. 2(b). Note that in this case, $X$ and $Y$ contain the same events, and therefore statistical correlations are present in the covariance matrix. The MicroBooNE Monte Carlo (MC) prediction before constraint is shown in red and the prediction after constraint is shown in blue. The shift between the red and blue predictions is dictated by the observation in data for the muon kinematics and by the model’s predicted relationship between the muon kinematics and the reconstructed hadronic energy. This difference between the red and blue predictions indicates how large of a correction to the central value is induced by the constraint through the rightmost term in Eq. 10. The bands surrounding the prediction, which show the uncertainties before and after constraint, illustrate the large reduction in uncertainties and enhanced sensitivity to mis-modeling that is obtainable with the conditional constraint. In particular, this test is expected to provide sensitivity to the modeling of the missing hadronic energy as is described in more detail in Sec. III.2. Other examples of similar constraint tests can be found in [10, 11, 12].
@@ III.2 Developing a Model Validation Procedure
However, this also brings about its own set of downsides. A combination of flux, detector and cross section effects could conspire in such a way to cancel out, thereby making reconstructed space tests less sensitive to mis-modeling that would still bias the extracted cross sections. The potential for such a situation is investigated in Sec. IV.2.1 and should be kept in mind when deciding which tests to include in the model validation procedure. In addition, it is important to note that the procedures we outline are more applicable when the background and signal models are both well defined and relatively well understood up to their uncertainties. If a signal significantly deviates from existing models or is completely unknown then these techniques cannot be properly applied as validation. In that case, other techniques should be employed to validate the background modeling and the signal efficiency within uncertainties before extracting a cross section.
@@
When developing a model validation procedure for a desired cross section extraction, one must make sure to design a set of tests that are sufficiently sensitive to mis-modeling in the phase space relevant for the cross section extraction. For example, validating only the overall event rate will not be sufficient for a differential cross section measurement. Such a test averages over the entire reconstructed phase space and will thus be insensitive to any mis-modeling related to the shape of the distribution that could bias the extraction of a differential cross section. Selecting an appropriately sensitive set of tests represents the key to a well-designed model validation procedure. To achieve this, one should consider what forms of mis-modeling may be capable of introducing bias into their cross section result and choose tests that address these possibilities. mis-modeling in distributions irrelevant to the extraction do not need to be examined. It may be useful to use FDSs to evaluate the ability of the selected set of tests to detect relevant mis-modeling before it begins to bias the extracted cross section results beyond stated uncertainties. A case study containing this type of FDS is presented in Sec. IV. This study mirrors the analysis done in [10] and thus also provides an example of what a full suite of tests used in a data-driven validation may look like. However, such studies are not mathematical proofs that the intended validation procedure constructed for the given analysis is guaranteed to detect relevant mis-modeling. As such, it is pertinent to probe the model from a variety of different angles in order to maximize the probability of detecting problematic forms of mis-modeling that would bias the cross section results beyond their uncertainties.
@@
The more challenging a variable or channel is to reconstruct or model, the more stringently it should be examined in the validation. For visible kinematic variables in more inclusive cross section measurements, a direct data to prediction comparison over the reconstructed distribution used in the unfolding is likely sufficient. However, more challenging distributions to model or reconstruct require additional validation. Examples include distributions containing events that are not fully contained within the detector volume, distributions used to map to quantities that cannot be fully reconstructed, or a more exclusive channel with kinematics cuts that are near detector thresholds. When selecting the constraining and constrained distributions, it is useful to keep in mind Eq. 10 and Eq. 11. The second term on the right of both these equations contain factors which dictate that smaller uncertainties on the constraining channel and larger correlations between the constraining and constrained channel will increase the power of the constraint.
@@
One particularly prominent example of this is the reconstruction of low energy protons. Low energy nucleons are known to be modeled significantly differently among event generators [43, 44] thus leading to very different predicted selection efficiencies for protons near the detection threshold [18]. This can have a large impact on measurements related to the proton’s energy, such as the kinetic energy of the leading proton or transverse kinematic imbalance variables [45, 46, 47], which are projections of the lepton and proton momentum onto the plane perpendicular to the neutrino direction. In such a case, a conditional constraint from a related visible kinematic variable less susceptible to the threshold may be useful. Examples of this include the energy and angle of the outgoing lepton. These constraints will reduce the allowed parameter space through correlations due to shared detector, flux, and cross section modeling thereby allowing for a more thorough investigation of the modeling of the proton kinematics near the detection threshold. Including an additional constraint from a closely related channel, such as one without protons, can provide additional information on the modeling of backgrounds or signal events that did not pass selection cuts and may also be useful in probing for relevant mis-modeling. This can be seen as similar to studying side-bands, which are often used to evaluate the modeling of backgrounds, but intends to evaluate the modeling used in the cross section extraction more directly. The central role of the correlations in this procedure makes it straightforward to evaluate if the model can consistently treat multiple distributions which provides a means of identifying when an observation in a side-band distribution is inconsistent with the observation in the signal distribution. This is a distinct advantage over evaluating side-bands in isolation, which does not provide such sensitivity. Though one could also obtain much of the same information through other approaches, such as tuning a model to a side-band channel or alternative kinematic variable, the ease with which one can rigorously account for and evaluate correlations between distributions makes the conditional constraint an attractive option.
@@
In the rightmost equality any contributions from the initial state nucleus have been absorbed into $E_{\mathrm{had}}^{\mathrm{missing}}$, that is, $E_{\mathrm{had}}^{\mathrm{missing}}$ is defined as the difference between the initial state neutrino energy and the observable leptonic and hadronic energy present in the final state. Through the use of $E_{\ell}^{\mathrm{rec}}$ as a constraint, the simultaneous measurement of $E_{\mathrm{had}}^{\mathrm{rec}}$ and $E_{\ell}^{\mathrm{rec}}$ is able to validate the predicted relationship between these distributions. This is achieved by using $E_{\ell}^{\mathrm{rec}}$ as channel $Y$ and $E_{\mathrm{had}}^{\mathrm{rec}}$ as channel $X$ in Eq. 10:
@@
where $\Sigma^{E_{\ell}^{\mathrm{rec}}E_{\ell}^{\mathrm{rec}}}$ describes the uncertainties on $E_{\ell}^{\mathrm{rec}}$ and $\Sigma^{E_{\ell}^{\mathrm{rec}}E_{\mathrm{had}}^{\mathrm{rec}}}$ describes the correlations between $E_{\ell}^{\mathrm{rec}}$ and $E_{\mathrm{had}}^{\mathrm{rec}}$. The same is done for Eq. 11 to obtain the reduced uncertainty band, $\Sigma^{E_{\mathrm{had}}^{\mathrm{rec}}E_{\mathrm{had}}^{\mathrm{rec}},\text{const.}}$, for the $E_{\mathrm{had}}^{\mathrm{rec}}$ prediction.
From Eq. 13, we see that the modeling of $E_{\ell}^{\mathrm{rec}}$, $E_{\mathrm{had}}^{\mathrm{rec}}$, and the correlations between them all play a role in obtaining the posterior prediction. Following Eq. 4, these can then be simultaneously examined with a GoF test on the constrained $E_{\mathrm{had}}^{\mathrm{rec}}$ prediction: